OpenStudy (anonymous):

y=2x-3/x+4 What is the equation of this functions vertical asymptote?

OpenStudy (psymon):

You find vertical asymptotes by finding the values of x that make the function undefined. So for this, the function will be undefined whenever the denominator is 0. SO that being said, we need to take the denominator (x+4), set it equal to 0 and then solve for x. Once you do that, your value of x will be where you have a vertical asymptote.

OpenStudy (anonymous):

your fast

OpenStudy (anonymous):

x=-4

OpenStudy (psymon):

Lol, Ididn't even notice this was your question x_x I just saw it and came here, haha. And yes, x= -4 is whereyour asymptote is :3

OpenStudy (anonymous):

funny! does this function have any additional asymptote like horizontal or oblique?

OpenStudy (anonymous):

horizontal... y=2

OpenStudy (anonymous):

is it 3/2

OpenStudy (psymon):

Horizontal yes. You have a horizontal asymptote when the degree of the numerator and denominator are the same. In this case, the highest power of x is 1 in the top and bottom, so your horizontal asymptote is dividing the coefficients of the highest power of x. As pg just posted, it would be y =2, since you have 2x divided by x.

OpenStudy (anonymous):

ow thanks @pgpilot326

OpenStudy (anonymous):

@Psymon when x has the same power right.

OpenStudy (anonymous):

yes, but you'll also have a horizontal asymptote at y=0 when the degree of the numerator is less than the degree of the denominator... just like y = 1/x

OpenStudy (psymon):

Well, when the highest power in the numerator and the denominator are the same you have a horizontal asymptote that you can find by dividing thecoefficients of those powers. And yeah, I always got someone watching over me to keep correcting my mistakes. You would also have one at y = 0 if the denominator is of a higher degree than the numeraotr.

OpenStudy (anonymous):

@pgpilot326 how did u get y=0

OpenStudy (psymon):

Sorry, I'll back off, pg has this better than I do x_x

OpenStudy (anonymous):

as x gets bigger , y = 1/x gets closer and closer to 0.

OpenStudy (anonymous):

no... you've got @Psymon

OpenStudy (anonymous):

i'll split.

OpenStudy (anonymous):

no @Psymon your helpful to you give good explanations too.

OpenStudy (psymon):

Lol, I just lack confidence and figure if anyone will be wrong about something it'll be me xD

OpenStudy (anonymous):

@Psymon and @pgpilot326 none of you guys have been wrong and you guys can team up.

OpenStudy (psymon):

Of course we can, I'm just paranoid xD

OpenStudy (anonymous):

no... trust me. you're doing great. some people here expect or want perfection but it's difficult to be that. I got pounced on for making mistakes but unfortunately it's part of the process. however the pouncing part shouldn't be... after all, I would hope everyone is here to learn.

OpenStudy (anonymous):

@Psymon your being really silly there is no reason to be paranoid.

OpenStudy (anonymous):

so please, @Psymon continue...

OpenStudy (psymon):

Lol, you don't need to explain that to me :P I'm just that way, I know I'm irrational in my thinking, haha.

OpenStudy (anonymous):

ya both of you guys continue.

OpenStudy (anonymous):

i just don't want to make anyone feel like others made me feel earlier today.

OpenStudy (psymon):

Well.....we finished up the horizontal asymptotes. As for the last kind of asymptote, oblique asymptotes, this would occur when the numerator has a degree exactly one higher than the denominator. If this is the case, then you find the equation for the oblique asymptote via long division. You divide the denominator into the numerator and ignore any remainder. What you come out with is your equation for the oblique asymptote :3 And sucks to here, pg D: I've left here for the day not always feeling too great myself x_x

OpenStudy (anonymous):

Dont worry about it even that happen to me and all i care about is if they got the help they needed.

OpenStudy (anonymous):

it's okay... good job on finishing up. Couldn't have said it any better! Keep up the good work and did you follow all of that @Rosa34 ?

OpenStudy (psymon):

I'm just more sensitive than most, just something I have to deal with xD But yeah, I know the situation didn't come up, but values of x that make the denominator 0 can be holes instead of asymptotes at times, too o.o

OpenStudy (anonymous):

I really think your smart @Psymon and @pgpilot326 just take the positive with you at the end of the day.

OpenStudy (anonymous):

thanks @Rosa34 , you too!

OpenStudy (anonymous):

your welcome

OpenStudy (psymon):

^_^

OpenStudy (anonymous):

That's right @Psymon! I think that happens when the factor(s) cancel.

OpenStudy (anonymous):

sorry got lost.

OpenStudy (anonymous):

you get the missing hole vs. the asymptote

OpenStudy (psymon):

Right. Feel like it's important to mention that. So yeah, you can sometimes have holes in your graph instead of asymptotes @Rosa34

OpenStudy (psymon):

Here's an example:

OpenStudy (anonymous):

suppose y = (x^2-1)/(x-1) this is just x+1 with a hole at x = 1.

OpenStudy (anonymous):

ok i think i understand

OpenStudy (anonymous):

sorry... didn't see cause i was typing. @Psymon probably has a better example and pic(?)

OpenStudy (psymon):

\[\frac{ x ^{2}+5x + 6 }{ x+2 }\] Now I would want to factor the top first before I start making any assumptions because if the denominator cancels, I have a hole instead of an asymptote \[\frac{ (x+3)(x+2) }{ (x+2) }\] Now -2 would make the denominator 0, so we would be undefined, but since that factor cancels out with the (x+2) factor on top, the -2 is a hole instead of asymptote. So basically you need to check and see what in the bottom cancels. If it cancels out, it's a hole, if it does not cancel out, it is a horizontal asymptote.

OpenStudy (anonymous):

very well said!!!

OpenStudy (psymon):

*vertical

OpenStudy (anonymous):

yes i got it thanks for both examples

OpenStudy (psymon):

Alrighty, awesome :3

OpenStudy (anonymous):

ok to sum it up yes there is a horizontal asymptote at y=2 and y=0.

OpenStudy (psymon):

No, not both.

OpenStudy (anonymous):

only at y=2

OpenStudy (psymon):

Correct :3

OpenStudy (anonymous):

and there is no oblique asymptotes

OpenStudy (psymon):

Correct. THat would only occur if the degree of the numerator was exactly one higher than the denominator.

OpenStudy (anonymous):

well, no. it's the line itself.

OpenStudy (anonymous):

in response to @Rosa34

OpenStudy (anonymous):

got it and to get y=2 you divide numorator and dinaminator which is 2/1

OpenStudy (psymon):

Well, you divide the coefficients of the highest powers.

OpenStudy (anonymous):

yeah as x goes to + or - infinity.

OpenStudy (anonymous):

like @Psymon said

OpenStudy (anonymous):

ok i understand

OpenStudy (anonymous):

so if i were to graph my horizontal and vertical asymptotes would it be two lines or the same line?

OpenStudy (psymon):

Well, you would draw a dashed line straight up and down through x = -4 and a dashed line straight left and right at y = 2.

OpenStudy (anonymous):

thank u very much @Psymon and @pgpilot326 for all the help.

OpenStudy (psymon):

Mhm mhm ^_^

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

so @Psymon i still didnt finish the problem we were working on last nigh. I really dont understand all of it.

OpenStudy (psymon):

I don't recall what it was.

OpenStudy (anonymous):

it was about the vertices

OpenStudy (psymon):

Ah. I just forget which conic section it was. The hyperbola, right?

OpenStudy (anonymous):

yes u are correct again.

OpenStudy (psymon):

Alrighty. Well, I'm not positive what exactly the problem was, but I can try and go over the rules for hyperbolas.

OpenStudy (anonymous):

ok so we got c=5 because a= 9, b=16, you use a^2+b^2=c^2

OpenStudy (psymon):

Ah, so this was \[\frac{ y ^{2} }{ 9 }-\frac{ x ^{2} }{ 16 }=1\] And yes, hyperbola gives you foci by a^2 + b^2 = c^2

OpenStudy (anonymous):

\[(y+3)^2/9- (x-2)^2/16=1\]

OpenStudy (anonymous):

ok what about vertices?

OpenStudy (psymon):

Oh, okay, so the center was elsewhere xD So yeah,center at 2, -3. Since this has y^2 as the positive fraction, the verticies and foci are up and down from the center. The verticies are up and down a units from the center. Now, a^2 is always the number in the denominator of the positive fraction, which means a^2 is 9 and a is 3. So up 3 units and down 3 units from the center our are verticies.